For a set of 15 observations the mean came out to be 17.3. Later on checking it is
discovered that an observation 19.5 was incorrectly recorded whereas the correct value was
15.9. Calculate the correct mean from these information.
1
Expert's answer
2020-07-06T20:12:29-0400

The mean (17.3) is the sum of the sampled values divided by the number of observations (15).

Therefore, the total sum is


"17.3\cdot 15=259.5"

Substracting the wrong value and adding the correct value, we receive the correct total sum:


"259.5-19.5+15.9=255.9"

Hence, the correct mean is


"\frac{255.9}{15}=17.06"

Answer: 17.06